Use Lagrange’s mean value theorem to prove that
x-x^2/2<log(1+x)<x-x^2/2(1+x) if x>0
1
Expert's answer
2020-04-20T12:21:44-0400

According to Lagrange’s Theorem, there exists at least one point ‘c’ on the open interval (a,b), such as

"f'(c)=\frac{f(b)-f(a)}{b-a}"  

Let's prove that "\log(x+1)>x-\frac{x^2}{2}"

Let's denote "f=\log(x+1)+\frac{x^2}{2}"

Using Lagrange’s Mean Value Theorem for function f, will lead us to

"\exist c\in (0,b)" such as "f'(c)=\frac{f(b)-f(0)}{b-0}=\frac{\log(b+1)+\frac{b^2}{2}}{b}"

Therefore

"f'(c)=\frac{1}{c+1}+c=\frac{\log(b+1)+\frac{b^2}{2}}{b}"

"\frac{1}{c+1}+c=1+\frac{c^2}{1+c}>1"

Hence

"\frac{\log(b+1)+\frac{b^2}{2}}{b}>1"

"\log(b+1)+\frac{b^2}{2}>b"

So "\log(x+1)>x-\frac{x^2}{2}"


Let's prove that "\log(x+1)<x-\frac{x^2}{2(1+x)}"

Let's denote "f=\log(x+1)+\frac{x^2}{2(x+1)}"

Using Lagrange’s Mean Value Theorem for function f, will lead us to "\exist c\in (0,b)" such as

"\frac{\log(b+1)+\frac{b^2}{2(b+1)}}{b}=f'(c)=\frac{1}{c+1}+\frac{2c2(c+1)-2c^2}{4(1+c)^2}=1-\frac{2c^2}{4(1+c)^2}<1"

Hence "\frac{\log(b+1)+\frac{b^2}{2(b+1)}}{b}<1"

"\log(b+1)+\frac{b^2}{2(b+1)}<b"

So "\log(x+1)<x-\frac{x^2}{2(1+x)}"